OpenStudy (anonymous):

A fair coin is tossed four times. Find the probability of obtaining: two heads and two tails

OpenStudy (anonymous):

ummm well first find how many different ways you can get two heads and two tails tossing 4 times... can you do that?

OpenStudy (anonymous):

yes i came up with 6 (HHTT, HTTH, TTHH, HTHT, THTH, THHT)

OpenStudy (anonymous):

good... and how many total outcomes are there?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

6/16

OpenStudy (anonymous):

good again... and just divide the two and you got it... but instead of writing down the sample space and finding all the ones with two H's and two T's, do you know how to put it in the calculator?

OpenStudy (anonymous):

but what if the coin was tossed a 100 times then how can we find the answer?

OpenStudy (anonymous):

thats where you would need the calculator for sure... so if it was tossed 100 times, how many times do they want H and T

OpenStudy (anonymous):

say 30 heads and 70 tails

OpenStudy (anonymous):

okay no problem... so in the calculator you are going to need to find the nCr, which stands for Combinations.... you could use that by tying 100nCr30 into the calculator or just use the formula which is (100!)/30!*70!

OpenStudy (anonymous):

if you have a TI 85 or above, its probably got to push MATH first then scroll to PRB and then scroll down.

OpenStudy (anonymous):

so the best way to do it is by using a calculator

OpenStudy (anonymous):

I would, for example the problem you had...if you type 4nCr2, which mean 4 combo 2, it will give you 6.

OpenStudy (anonymous):

keep in mind though, this is only good for when order is not important

OpenStudy (anonymous):

there is this last question that is bothering me. i know p=3/5, n=5 but i dont know what k is equal to

OpenStudy (anonymous):

what is the entire question

OpenStudy (anonymous):

the question is: a weighted coin is biased such that a tail comes up 60% of the time. the coin tossed 5 times. find the probabability of obtaining: tails on the first four tosses only

OpenStudy (anonymous):

okay, so what is the probability that there is a tails if we were to only toss it once?

OpenStudy (anonymous):

actually more specific, the FIRST toss

OpenStudy (anonymous):

sorry i couldnt get back to you because something was wrong with the website

OpenStudy (anonymous):

no biggie. did you figure it out.

OpenStudy (anonymous):

no i couldnt but answering your question it should be 1/2 for each toss

OpenStudy (anonymous):

but the problem is that it is biased 60/100=3/5

OpenStudy (anonymous):

are the the four tosses treated as independently?

OpenStudy (anonymous):

so the first toss probability of tails is (3/5) second (3/5) third (3/5) fourth (3/5) but finall on the last toss you want heads so that is 2/5. so to find probability you just take 3/5 four times multiplied by 2/5 p=(3/5)^4(2/5)

OpenStudy (anonymous):

so should be around .05

OpenStudy (anonymous):

the answer is correct :)

OpenStudy (anonymous):

so when coins are biased we cant use the formula to find the answer?

OpenStudy (anonymous):

what formula do they give you?

OpenStudy (anonymous):

bernolli

OpenStudy (anonymous):

you could use that formula if order wasnt important... but the fact that they tell you that you must get a tails the first four times no longer makes this problem a combination problem... if it would of said you toss it 5 times and all they want is just 4 tails, then you could, b/c that will let you get a heads on any of the 5 tosses as long as it happens only once... your problem says the heads must occur on the final toss

OpenStudy (anonymous):

thanks a lot. i really appreciate your help :) and i really like how you explained it to me

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

could you just have a look if i am answering this question right?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

55% of the local municipality support the local council. If 8 people are selected at random, find the probability that: half support the council

OpenStudy (anonymous):

k=8, p=55/100=11/20

OpenStudy (anonymous):

am i right here?

OpenStudy (anonymous):

no, give me a few. should be more than that i would think.

OpenStudy (anonymous):

mmm i am clueless :)

OpenStudy (anonymous):

do they want exactly 4 or atleast 4?

OpenStudy (anonymous):

if 8 people are selected randomly

OpenStudy (anonymous):

right, 8 people are chosen, but to me it is asking the probability of 4 supporting the council.... OR , atleast 4 support the council

OpenStudy (anonymous):

would it be n=8 and k=4

OpenStudy (anonymous):

no it is asking for the probability of 4 supporting the council

OpenStudy (anonymous):

if the problem is asking for 4 exactly out of the 8, then use the bernoulli formula. 8nCr4*.55^4*.45^4

OpenStudy (anonymous):

ill have a look if it is the right answer. it looks right to me

OpenStudy (anonymous):

it should be .263 or so

OpenStudy (anonymous):

your right as always :)

OpenStudy (anonymous):

did you understand how we did that?

OpenStudy (anonymous):

yes i got it so n=8 and k=4

OpenStudy (anonymous):

if you have one that says atleast 4, then that would be a total different calculation... but let me know if you got any more.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

The probability of Colin beating Maria at golf is 0.4. If they play once a week throughout the entire year and the outcome of each game is independent of any other, find the probability that they will have won the same number of matches

OpenStudy (anonymous):

I guess p=0.4 and q=0.6

OpenStudy (anonymous):

Would n be equal to the number of games they played?

OpenStudy (anonymous):

And k will be equal to half of n?

OpenStudy (anonymous):

yeah same formula (52nCr26)(.4^26)(.6^26) = approx. .038

OpenStudy (anonymous):

How can we calculate the factorial. Is there a factorial calculator that can give me the answer

OpenStudy (anonymous):

TI calculators will do it... anyone that is above a TI-83 i believe.

OpenStudy (anonymous):

A new drug being trialled gives 8% of the patients a violent reaction. If 200 patients trial the drug, find the probability that 12 patients have a violent reaction to the drug

OpenStudy (anonymous):

P=0.08

OpenStudy (anonymous):

Will n be 20?

OpenStudy (anonymous):

Sorry 200

OpenStudy (anonymous):

And k =12

OpenStudy (anonymous):

@zonazoo